OpenStudy (anonymous):

What are the solutions of x(x - 2) = 5?

OpenStudy (anonymous):

distribute out to find x^2 - 2x - 5 = 0, solve

OpenStudy (anonymous):

use the quadratic formula

OpenStudy (anonymous):

I forget formula

OpenStudy (anonymous):

just used quadratic formula to solve for x: a=1 b=-2 and c=-5

OpenStudy (anonymous):

yes what is the quadrati formula

OpenStudy (anonymous):

please tell me

OpenStudy (anonymous):

Difficult to type: http://en.wikipedia.org/wiki/Quadratic_equation#Quadratic_formula

OpenStudy (anonymous):

where a, b, and c come from the quadratic expression you have (ax^2 + bx + c = 0)

OpenStudy (anonymous):

ok I saw it now thanks

OpenStudy (anonymous):

\[x=\frac{-b\pm\sqrt{b^2 -4ac}}{2a}\] Type x=\frac{-b\pm\sqrt{b^2 -4ac}}{2a} in equation editor

OpenStudy (anonymous):

ok I got it

OpenStudy (anonymous):

so what is the result?

OpenStudy (anonymous):

my result is not right

OpenStudy (anonymous):

-1 plus or minus sqrt 6 Thought you would have worked out yourself by now...

OpenStudy (anonymous):

-2 plus minus sqrt(4 +20)/2

OpenStudy (anonymous):

x= 2+ squared 4- 4.5/2

OpenStudy (anonymous):

but b= minus2

OpenStudy (anonymous):

u need to be more careful with your signs b =-2 and it's 4*-5

OpenStudy (anonymous):

their answer is 1 - sqrt(6) and 1 + sqrt(6)

OpenStudy (anonymous):

why sqrt(6)?

OpenStudy (anonymous):

sqrt(24) = sqrt(4)sqrt(6) and the sqrt(4) = 2 cancels with the 2 on the bottom.

myininaya (myininaya):

x(x-2)=5 x^2-2x=5 x^2-2x+1=5+1 (x-1)^2+6 x-1= plus or minus sqrt{6} x=1 plus or minus sqrt{6} x=1+sqrt{6} or x=1-sqrt{6}

OpenStudy (anonymous):

oh I see thank you

myininaya (myininaya):

oops that one line should read (x-1)^2=6 not (x-1)^2+6 lol

OpenStudy (anonymous):

yes